How do I simplify the rational expression? (book says)you will need to use factoring by grouping.
ax-3x+2ay-6y
____________
2ax-6x+ay-3y

Answer by mananth(16949)

ax-3x+2ay-6y
____________
2ax-6x+ay-3y
=x(a-3)+2y(a-3) / 2x(a-3)+y(a-3)
=(x+2y)(a-3) / (2x+y)(a-3)
=(x+2y) /(2x+y)